# Adaptive Grid-Based Thompson Sampling for Efficient Trajectory Discovery

### v1

@@ 2.5 Bayesian Optimization with Thompson Sampling
Bayesian optimization is a sequential design framework for optimizing expensive black-box functions (garnett2023bayesian, ; Jones1998, ). It proceeds by constructing a probabilistic surrogate model, such as a GP over the objective function, and uses posterior uncertainty to guide the selection of future evaluation points.
Among acquisition strategies, Thompson Sampling (TS) (thompson1933likelihood, ) is a simple yet effective approach that draws a sample from the posterior distribution of the surrogate model and selects the input corresponding to the optimum of the sampled function. This randomized strategy implicitly balances exploration and exploitation. In each iteration, a surrogate $f\sim\mathcal{GP}(\mu,k)$ is sampled over a candidate grid, and the next input is chosen as the optimizer of the sampled realization.
@@ 3 Thompson Sampling Over An Adaptive Grid
To support scalable BO for efficient trajectory finding, we adopt a discrete search strategy wherein TS is performed over a finite set of candidate input points. This grid-based formulation, as outlined in Algorithm 1, offers a computationally tractable alternative to continuous optimization, especially in the context of simulation-based models where each function evaluation can be costly. By limiting evaluations to a fixed set of locations, the approach avoids expensive numerical optimization routines (agrawal2013thompson, ) and enables efficient batch evaluations in HPC settings (bergstra2012random, ; gonzalez2016batch, ).
Grid-based TS has been widely used in multi-armed bandit problems (chapelle2011empirical, ), hyperparameter tuning (snoek2012practical, ), and the calibration of stochastic simulators (frazier2018tutorial, ; fadikar2023TOO, ). While the fixed-grid approach simplifies implementation and maintains statistical efficiency, its performance depends heavily on the resolution and coverage of the search space. Coarse grids may miss important regions of the parameter space, particularly in high-dimensional or nonlinear settings. Recent developments address this limitation through adaptive discretization schemes. For instance, binois2020kalai propose iteratively refining the grid to enhance coverage, while gramacy2022triangulation use geometric structures to select informative candidate points. These strategies retain the computational advantages of grid-based search while enabling greater flexibility and improved optimization performance. In the next section, we describe our adaptive grid strategy tailored for trajectory-level optimization using CRNGP surrogates.
@@ 3.1 Adaptive Grid Refinement Strategy for Efficient Thompson Sampling
To address the limitations of searching over a fixed-grid, we introduce an adaptive grid refinement strategy that enhances TS performance by dynamically updating the search space at each iteration. This approach aims to focus computational effort on regions with high potential for matching observed data while avoiding unnecessary evaluations in less informative areas. This idea is conceptually related to strategies such as trust-region modeling and space partitioning, both of which adaptively localize the search. Trust-Region Bayesian Optimization (TuRBO) (eriksson2019scalable, ) confines the sampling to high-performing neighborhoods, whereas space-partitioning methods (bubeck2011bandit, ; wang2020lamcts, ) divide the parameter space and allocate more resources to promising subregions. Building on these ideas, we implement an intuitive and scalable alternative. Specifically, we design a two-stage refinement mechanism applied to the candidate grid $\tilde{\mathbf{X}}_{t}$ at each BO iteration $t$:
@@ 3.1.2 Densification Step
Step 5 of Algorithm 1 requires sampling from the GP posterior on the $M$ candidates. For large $M$, it may become computationally demanding, with a $\mathcal{O}(M^{3})$ complexity. Assuming that all seeds are considered across the unique $\mathbf{x}$’s among the $M$ candidates, the resulting virtual candidate set with $\tilde{M}$ candidates leads to a covariance matrix over all possible ($\mathbf{x},r$) combinations. Given the separable covariance structure between the $\mathbf{x}$’s and $r$’s, this covariance matrix has a Kronecker structure, leading to a faster Cholesky decomposition of the covariance over all $\mathbf{x}$’s and over all $r$’s separately. Then, based on the technique described by chiles2012geostatistics ; chevalier2015fast , these unconditional simulations can be conditioned by subtracting the GP prediction on this realization before adding the prediction on the actual data.
@@ 4.1 Model
We generate synthetic outbreaks with a discrete–time stochastic SIR model implemented using the odin R package fitzjohn2024odin . In a typical SIR model, a population is partitioned into compartments, and the following system of ordinary differential equations describes how individuals transition between compartments:
@@ 4.2 Experiments
These experiments were performed on Argonne Laboratory Computing Resource Center’s Improv cluster using the EMEWS framework (ozikDesktopLargeScaleModel2016a, ; collierDistributedModelExploration2024, ). All BO algorithms were implemented in R. All the code used in this paper is publicly available at https://github.com/emews/adaptiveTS.
Table 2: Comparator approaches for TS-based BO
@@ 4.3 Results
To evaluate both the speed and overall utility of each method in surfacing high-quality model trajectories, we approach our analysis through a thresholding lens. Figure 2 shows all the trajectories explored by aCRN for a particular experiment configuration, colored by a set of RMSE threshold categories with respect to a simulated ground truth (black dots). We note that acceptable thresholds will vary based on stakeholder needs and use-case.
Figure 2: Simulated epidemic trajectories generated by the aCRN method for a single experiment configuration. Each trajectory is colored according to its RMSE category, with thresholds at 15, 20, 25, and 30.
@@ 4.3.1 Collecting high-quality trajectories
We also compare results from the best-performing experimental design for each method. Specifically, we use the median proportion of below-threshold trajectories across the 10 experimental replicates for each set of hyper-parameters. We observe that here too aCRN considerably outperforms the comparators (see Figure 8 in Supplementary Materials). We also observe that sequential algorithms benefit from smaller initial budgets (see Figure 10).
@@ 4.3.2 Fast time-to-solution
In addition to assessing the overall quality of the trajectories each method identifies within a given simulation budget, we want to assess the time-to-solution of each method–that is, how quickly high-quality trajectories are found through the optimization procedure. The time-to-solution is measured by counting the number of trajectories with RMSE below the specified threshold per budget expenditure. To quantify this metric across experiments, we estimate the area under the curve (AUC) of each search pathway using the standard trapezoidal formulation, and report a relative AUC (rAUC), normalized by $(N_{max})^{2}$. Specifically, letting $QT_{t}$ be the number of quality trajectories found by time $t$,
$\text{rAUC}:=\frac{1}{(N_{max})^{2}}\sum_{t=2}^{N_{max}}\frac{(QT_{t}+QT_{t-1})}{2}$ (12)
This metric can highlight differences between methods that are not captured by the previous metric. Figure 4 shows two example trajectory-finding pathways from different experiments using aCRN (configuration 1) and fHet (configuration 2); we see that the method that collects the most below-threshold trajectories at the end of the simulation budget has a lower rAUC than the other experiment, which collects high-quality trajectories at a faster rate.
Figure 4: Number of trajectories with RSME ¡ 15 found using aCRN and fHet for two different experiments: Configuration 1 = $(n_{init}:10,n_{rep}:10,M:300,n_{TS}:10,\text{method: aCRN})$, Configuration 2 = $(n_{init}:5,n_{rep}:10,M:300,n_{TS}:20,\text{method: fHet})$. The rAUC values for each experiment are shown at the end of the search pathways.
In the Supplementary Materials we show the distribution of this metric across all experiments, where we again observe that aCRN consistently achieves the highest rAUC across simulation budgets and thresholds (Figure 9). We also include additional experiments to demonstrate the robustness of the results.
@@ 4.3.3 Parameter exploration
Figure 5 compares the sampling behavior of the two top performing methods, fHet and aCRN, across the joint model parameter space $(\beta,\gamma)$ and replicate identifiers $r$. Each red dot represents an evaluated input-replicate pair $(\mathbf{x},r)$, and the shaded plane indicates the restricted replicate set used for each. The black dot denotes the ground-truth parameter setting used to generate the observed data.
Figure 5: Sampling behavior of the two top performing methods, fHet and aCRN, across the joint model parameter space $(\beta,\gamma)$ and replicates. Panels a) and b) show the samples for fHet. Panels c) and d) show the samples for aCRN. The black dots in panels a) and c) indicate the location of the ground-truth parameter setting used in the experiments. Panels b) and d) show the density of samples in the two dimensional $(\beta,\gamma)$ space.
We observe that fHet exhibits broad sampling along the replicate axis (Figure 5a)) but limited diversity across the parameter space (Figure 5b)). In contrast, aCRN, which explicitly includes the replicates in its search space, achieves broader parameter coverage (Figure 5c) and d)).
@@ 5 Finding Trajectories in a City-scale Agent-based Model
The stochastic SIR model in the preceding section provides a convenient testbed for evaluating optimization strategies, but it is too simplistic to capture many of the complexities of real-world epidemics. Here we illustrate the performance of our proposed method on a more realistic and computationally expensive model by applying it to CityCOVID (Ozik2021, ; macal2020citycovid, ), a high-resolution ABM developed during the COVID-19 pandemic to support public health decision making in the Chicago metropolitan area.
@@ 5.1 CityCOVID simulation model
Historically, calibrating such a complex model has posed considerable challenges. The initial calibration of CityCOVID used incremental mixture approximate Bayesian computation (IMABC) (rutter2019microsimulation, ), a likelihood-free method that required approximately 35,000 simulations to arrive at a plausible posterior distribution. While recent efforts have attempted to reduce the computational burden and improve uncertainty quantification (e.g., via surrogate models (Ozik2021, ; binoisPortfolioApproachMassively2025, ; robertson2025bayesian, )), these approaches have focused on matching the mean response of the simulator to observed data. Such strategies may obscure important trajectory-level variations and fail to fully account for the structural role of randomness in shaping epidemic dynamics. As shown in Figure 1, even parameter configurations calibrated to the mean response can yield a wide range of individual trajectories, many of which deviate significantly from the observed data, while their mean behavior is consistent with observed data.
Initial CityCOVID calibration efforts (Ozik2021, ) identified nine relevant parameters via a Morris global sensitivity analysis (morris1991, ), which was shown to be further reducible to four parameters in robertson2025bayesian . Here we further constrain one of these parameters, the time of initial exposure to infection, to a fixed value and focus on three parameters (Table 3).

### v2

@@ 2.5 Bayesian Optimization with Thompson Sampling
Bayesian optimization is a sequential design framework for optimizing expensive black-box functions (Garnett, 2023; Jones et al., 1998). It proceeds by constructing a probabilistic surrogate model, such as a GP over the objective function, and uses posterior uncertainty to guide the selection of future evaluation points.
Among acquisition strategies, Thompson Sampling (TS) (Thompson, 1933) is a simple yet effective approach that draws a sample from the posterior distribution of the surrogate model and selects the input corresponding to the optimum of the sampled function. This randomized strategy implicitly balances exploration and exploitation. In each iteration, a surrogate $f\sim\mathcal{GP}(\mu,k)$ is sampled over a candidate grid, and the next input is chosen as the optimizer of the sampled realization.
@@ 3 Thompson Sampling Over An Adaptive Grid
To support scalable BO for efficient trajectory finding, we adopt a discrete search strategy wherein TS is performed over a finite set of candidate input points. This grid-based formulation, as outlined in Algorithm 1, offers a computationally tractable alternative to continuous optimization, especially in the context of simulation-based models where each function evaluation can be costly. By limiting evaluations to a fixed set of locations, the approach avoids expensive numerical optimization routines Agrawal and Goyal (2013) and enables efficient batch evaluations in HPC settings Bergstra and Bengio (2012); Gonzalez et al. (2016).
Grid-based TS has been widely used in multi-armed bandit problems Chapelle and Li (2011), hyperparameter tuning Snoek et al. (2012), and the calibration of stochastic simulators Frazier (2018); Fadikar et al. (2023). While the fixed-grid approach simplifies implementation and maintains statistical efficiency, its performance depends heavily on the resolution and coverage of the search space. Coarse grids may miss important regions of the parameter space, particularly in high-dimensional or nonlinear settings. Recent developments address this limitation through adaptive discretization schemes. For instance, Binois et al. (2020) propose iteratively refining the grid to enhance coverage, while Gramacy et al. (2022) use geometric structures to select informative candidate points. These strategies retain the computational advantages of grid-based search while enabling greater flexibility and improved optimization performance. In the next section, we describe our adaptive grid strategy tailored for trajectory-level optimization using CRNGP surrogates.
@@ 3.1 Adaptive Grid Refinement Strategy for Efficient Thompson Sampling
To address the limitations of searching over a fixed-grid, we introduce an adaptive grid refinement strategy that enhances TS performance by dynamically updating the search space at each iteration. This approach aims to focus computational effort on regions with high potential for matching observed data while avoiding unnecessary evaluations in less informative areas. This idea is conceptually related to strategies such as trust-region modeling and space partitioning, both of which adaptively localize the search. Trust-Region Bayesian Optimization (TuRBO) Eriksson et al. (2019) confines the sampling to high-performing neighborhoods, whereas space-partitioning methods Bubeck and Cesa-Bianchi (2011); Wang and He (2020) divide the parameter space and allocate more resources to promising subregions. Building on these ideas, we implement an intuitive and scalable alternative. Specifically, we design a two-stage refinement mechanism applied to the candidate grid $\tilde{\mathbf{X}}_{t}$ at each BO iteration $t$:
@@ 3.1.2 Densification Step
Figure 3 illustrates how adaptive grid TS combined with a CRNGP surrogate refines the search for parameter-seed tuples $(X,r)$ that minimize the discrepancy $Y=d(f(X,r),\mathbf{y}_{\text{obs}})$. The example is based on the stochastic SIR model described in section 4, and is constructed in a one-dimensional setting by fixing the recovery rate at $\gamma=0.35$ (normalized to the $[0,1]$ scale) and varying the transmission rate $X=\beta$. Three stochastic replicates are considered, corresponding to distinct random seeds. Synthetic ground truth observation is generated by setting $\beta=0.45,\;r=5$, and the objective of the optimization is to identify pairs $(X,r)$ for which the simulated trajectory closely matches the observed data, i.e., to have minimum $Y$ values. The discrepancy values are log-transformed and standardized to have mean zero and unit variance prior to fitting the surrogate and applying BO.
The left panels show the evolution of the candidate grid across successive BO iterations. Vertical colored lines represent the set of candidate input locations considered at each iteration, and the plotting symbols indicate the subset of locations that are evaluated. As the iterations progress, the grid increasingly concentrates around regions associated with lower predicted discrepancy values as a result of likelihood-based filtering and grid densification. The right panels display the corresponding CRNGP posterior summaries for each replicate, where solid lines denote posterior means and shaded regions indicate posterior uncertainty. As additional evaluations are incorporated, uncertainty decreases in the region of interest (near $X=0.45)$. This simple 1-D example demonstrates how the CRNGP surrogate enables replicate-aware inference, and how adaptive grid refinement guides exploration toward regions of the augmented input space that are most likely to yield high-quality trajectories.
@@
Figure 3: Illustration of adaptive grid TS with a CRNGP surrogate to find $(X,r)$ tuple having minimal $Y$ values. The example consists of three replicates denoted by different colors. Left panels show the evolving grid locations in the augmented input space, with colored vertical lines indicating candidate input locations (and replicate) at each BO iteration. Among those candidate locations, the evaluated points are indicated by corresponding plotting characters. Right panels show the CRNGP posterior mean (solid lines) and uncertainty bands for each replicate using the evaluated points until the corresponding iteration. The red dashed vertical line indicates the parameter value at which the ground-truth is generated. From top to bottom, rows correspond to increasing BO iterations.
@@
Step 5 of Algorithm 1 requires sampling from the GP posterior on the $M$ candidates. For large $M$, it may become computationally demanding, with a $\mathcal{O}(M^{3})$ complexity. Assuming that all seeds are considered across the unique $\mathbf{x}$’s among the $M$ candidates, the resulting virtual candidate set with $\tilde{M}$ candidates leads to a covariance matrix over all possible ($\mathbf{x},r$) combinations. Given the separable covariance structure between the $\mathbf{x}$’s and $r$’s, this covariance matrix has a Kronecker structure, leading to a faster Cholesky decomposition of the covariance over all $\mathbf{x}$’s and over all $r$’s separately. Then, based on the technique described by Chiles and Delfiner (2012); Chevalier et al. (2015), these unconditional simulations can be conditioned by subtracting the GP prediction on this realization before adding the prediction on the actual data.
@@ 4.1 Model
We generate synthetic outbreaks with a discrete–time stochastic SIR model implemented using the odin R package FitzJohn (2024). In a typical SIR model, a population is partitioned into compartments, and the following system of ordinary differential equations describes how individuals transition between compartments:
@@ 4.2 Experiments
Our experiments explore the conditions under which we can find trajectories sufficiently close to the ground-truth within a fixed simulation budget $N_{max}$. We initialize each optimization algorithm we evaluate with a space-filling design of $n_{init}$ model parameters, each of which is replicated $n_{rep}$ times for a total initial cost of $N_{0}=n_{init}\cdot n_{rep}$ simulation runs. Let $\mathbf{y}_{ij}=f_{T}(\beta_{i},\gamma_{i},r_{j})$ represent the $T$-dimensional trajectory output of the SIR model, evaluated at $(\beta_{i},\gamma_{i})$ with the random seed set to $r_{j}$. Given the ground-truth trajectory $\mathbf{y}_{\text{obs}}$, $z_{ij}:=\|\mathbf{y}_{ij}-\mathbf{y}_{\text{obs}}\|_{2}^{2}$ is the scalar output for training the GP surrogates on and the quantity we optimize over. After the GP surrogate is fit on the initial design, additional $(\beta,\gamma,r)$ locations are selected to evaluate through batched TS algorithms until a total of $N_{max}$ simulations are run. Each iteration of the TS procedure selects $n_{TS}$ new points to evaluate from a grid (either fixed or adaptive) of size $M$. Each of these hyperparameters ($N_{max}$, $n_{init}$, $n_{rep}$, $n_{TS}$, $M$) has the potential to impact the performance of the algorithms, and we perform experiments for all combinations (see Table 1), replicating each experiment 10 times to account for algorithmic stochasticity (note this is separate from specifying the simulation random seed) for a total of 990 experimental runs.
@@
Table 2: Comparator approaches for TS-based BO
These experiments were performed on Argonne Laboratory Computing Resource Center’s Improv cluster using the EMEWS framework Ozik et al. (2016); Collier et al. (2024). All BO algorithms were implemented in R. All the code used in this paper is publicly available at https://github.com/emews/adaptiveTS.
@@ 4.3 Results
To evaluate both the speed and overall utility of each method in finding high-quality model trajectories, we approach our analysis through a thresholding lens. Figure 4 shows all the trajectories explored by aCRN for a particular experiment configuration, colored by a set of RMSE threshold categories with respect to a simulated ground truth (black dots). We note that acceptable thresholds will vary based on stakeholder needs and use-case.
Figure 4: Simulated epidemic trajectories found by the aCRN method for a single experiment configuration. Each trajectory is colored according to its RMSE category, with thresholds at 15, 20, 25, and 30. The ground truth is shown using black dots.
@@ 4.3.1 Collecting High-Quality Trajectories
Figure 6: Means and standard errors of proportions of below-threshold trajectories found for the best-performing experimental design for each method.
Figure 7: Proportion of trajectories with RMSE below 30 across all experiments varying the hyper-parameters, stratified by simulation budget $N_{\max}$, and initial design size $N_{0}$.
We also compare results from the best-performing experimental design for each method. For this comparison, we compute, for each method, the median proportion of below-threshold trajectories across the 10 experimental replicates associated with its optimal hyperparameter setting. As shown in Figure 6, aCRN consistently finds a higher proportion of trajectories with RMSE below each threshold across all simulation budgets $N_{\max}\in\{300,500,700\}$. The performance gap between aCRN and the comparator methods is most notable at stricter thresholds and higher $N_{\max}$, indicating that aCRN is more efficient in concentrating evaluations in regions of the (parameter, seed) space that yield high-quality trajectories.
Furthermore, Figure 7 shows that all algorithms, being sequential in nature benefit from smaller initial designs, that is, allocating a larger fraction of the total simulation budget to adaptive sampling rather than to the initial space-filling design. This enables the acquisition mechanism to more efficiently sample the (parameter, seed) space over informative regions. Here too, we observe that aCRN outperforms the other methods in collecting more high-quality trajectories across different simulation budget and initial design size.
@@ 4.3.2 Fast Time-to-Solution
In addition to assessing the overall quality of the trajectories each method identifies within a given simulation budget, we want to assess the time-to-solution of each method, that is, how quickly high-quality trajectories are found through the optimization procedure. This metric is particularly relevant in epidemiology, where simulation runs can be expensive and timely inference is essential for updating forecasts or informing real-time decision-making. Methods that identify promising trajectories earlier in the search process allow resources to be allocated more efficiently and reduce the turnaround time required to obtain actionable outputs.
The time-to-solution is measured by counting the number of trajectories with RMSE below the specified threshold per budget expenditure. To quantify this metric across experiments, we estimate the area under the curve (AUC) of each search pathway using the standard trapezoidal formulation, and report a relative AUC (rAUC), normalized by $(N_{max})^{2}$. Specifically, letting $QT_{t}$ be the number of quality trajectories found by time $t$, we define
$\text{rAUC}:=\frac{1}{(N_{max})^{2}}\sum_{t=2}^{N_{max}}\frac{(QT_{t}+QT_{t-1})}{2},$ (12)
where higher values of rAUC correspond to methods that identify high-quality trajectories earlier and more consistently over time. This metric highlights temporal differences between methods that may not be captured by end-of-budget performance alone. Figure 8 illustrates two representative search pathways from different experiments using aCRN (configuration 1) and fHet (configuration 2). Although configuration 1 ends up with lower number of below-threshold trajectories at the end of full simulation budget of 700, it discovers high-quality trajectories much earlier in the search process, achieving a higher rAUC. This emphasizes that end-of-budget summaries may obscure meaningful differences in search efficiency.
Figure 8: Number of trajectories with RSME << 15 found using aCRN and fHet for two different experiments: Configuration 1 = $(n_{init}:10,n_{rep}:10,M:300,n_{TS}:10,\text{method: aCRN})$, Configuration 2 = $(n_{init}:5,n_{rep}:10,M:300,n_{TS}:20,\text{method: fHet})$. The rAUC values for each experiment are shown at the end of the search pathways.
In Figure 9 we show the distribution of this metric across all experiments, where we again observe that aCRN consistently achieves the highest rAUC across simulation budgets and thresholds. The results suggest that aCRN is not only able to find more high-quality trajectories, but also does so at a faster rate compared to other methods. We also include additional experiments in the supplementary materials to demonstrate the robustness of the results.
Figure 9: rAUC of best performing design for each method under different budget and varying thresholds.
@@ 4.3.3 Parameter Exploration
In the case of a stochastic epidemiological model, such as our simple SIR model, or a more complex one such as CityCOVID, calibration parameters typically interact with underlying model stochasticity, such as population mixing patterns. Since each random seed induces a different mixing pattern for the present SIR model, the best set of calibration parameters may differ from one seed to another. For example, a random seed leading to more contacts than another may identify a lower $\beta$ in the former and higher in the latter, since the infection rate models both inherent pathogen transmissibility as well as population contact structure. Thus, it is of interest to study whether a TOO approach can facilitate increased exploration of parameter space to identify suitable trajectories. Figure 10 compares the sampling behavior of the six methods across the joint model parameter space $(\beta,\gamma)$ and replicate identifiers $r$. Each red dot represents an evaluated input-replicate pair $(\mathbf{x},r)$, and the shaded plane indicates the restricted replicate set used for each. Figure 11 shows the densities of the samples projected to the ($\beta$$\gamma$) plane. The black dot denotes the ground-truth parameter setting used to generate the observed data.
Figure 10: Sampling behavior of the comparator methods along with the initial design, across the joint model parameter space $(\beta,\gamma)$ and replicates. The red circles show the samples from each of the methods. The black dots each each panel indicate the location of the ground-truth parameter setting used in the experiments.
Figure 11: Sampling behavior of all the methods along with the initial design, across the two dimensional model parameter $(\beta,\gamma)$, and the density of samples in the background.
We observe that fHet exhibits broad sampling along the replicate axis (Figure 10)) but limited diversity across the parameter space (Figure 11)). In contrast, aCRN, which explicitly includes the replicates in its search space, achieves broader parameter coverage (Figure 10c) and d)).
@@ 5 Finding Trajectories in a City-scale Agent-based Model
The stochastic SIR model in the preceding section provides a convenient testbed for evaluating optimization strategies, but it is too simplistic to capture many of the complexities of real-world epidemics. Here we illustrate the performance of our proposed method on a more realistic and computationally expensive model by applying it to CityCOVID Ozik et al. (2021); Macal et al. (2020), a high-resolution ABM developed during the COVID-19 pandemic to support public health decision making in the Chicago metropolitan area.
@@ 5.1 CityCOVID Simulation Model
Historically, calibrating such a complex model has posed considerable challenges. The initial calibration of CityCOVID used incremental mixture approximate Bayesian computation (IMABC) Rutter et al. (2019), a likelihood-free method that required approximately 35,000 simulations to arrive at a plausible posterior distribution. While recent efforts have attempted to reduce the computational burden and improve uncertainty quantification (e.g., via surrogate models Ozik et al. (2021); Binois et al. (2025); Robertson et al. (2025)), these approaches have focused on matching the mean response of the simulator to observed data. Such strategies may obscure important trajectory-level variations and fail to fully account for the structural role of randomness in shaping epidemic dynamics. As shown in Figure 2, even parameter configurations calibrated to the mean response can yield a wide range of individual trajectories, many of which deviate significantly from the observed data, while their mean behavior is consistent with observed data.
Initial CityCOVID calibration efforts Ozik et al. (2021) identified nine relevant parameters via a Morris global sensitivity analysis Morris (1991), which was shown to be further reducible to four parameters in Robertson et al. (2025). Here we further constrain one of these parameters, the time of initial exposure to infection, to a fixed value and focus on three parameters (Table 3).